OpenStudy (anonymous):

HELP PLEASE!!!! The average air pressure in another town is 29 inches but the actual pressure is within 4 inches of the average. Write and solve an inequality to find the range of air pressure in this town.

OpenStudy (anonymous):

@helplesspenguin

ganeshie8 (ganeshie8):

the difference between, 'actual' and '29' <= 4

ganeshie8 (ganeshie8):

|x - 29| <= 4 solve x

OpenStudy (anonymous):

x<=33???

ganeshie8 (ganeshie8):

thats upper limit, wat about lower limit... how low the pressure can go ?

OpenStudy (anonymous):

25?

ganeshie8 (ganeshie8):

yes so wats the range of values the pressure can take ?

OpenStudy (anonymous):

8..

ganeshie8 (ganeshie8):

x <= 33 and x >= 25

ganeshie8 (ganeshie8):

so the range is : 25 to 33

ganeshie8 (ganeshie8):

you can show it as :- \(25 \le x \le 33\)

OpenStudy (anonymous):

oh... so how would I write my answer?

ganeshie8 (ganeshie8):

^ above

OpenStudy (anonymous):

Ahhhh :) Thank you

ganeshie8 (ganeshie8):

np :)